# thanks: These two authors contributed equally

### v1

@@ Abstract
We model and investigate the collective nonlinear optical response of an ensemble of two-level emitters that are weakly coupled to a single-mode waveguide. Our approach generalizes the insight that photon-photon correlations in the light scattered by a single two-level emitter result from two-photon interference to the case of many emitters. Using our model, we study different configurations for probing the nonlinear response of the ensemble, e.g., through the waveguide or via external illumination, and derive analytical expressions for the second-order quantum coherence function, $g^{(2)}(\tau)$, as well as for the squeezing spectrum of the output light in the waveguide, $S_{\theta}(\omega)$. For the transmission of resonant guided light, we recover the same predictions as previously made with far more involved theoretical models when analyzing experimental results regarding $g^{(2)}(\tau)$ (Prasad et al. [1]) and $S_{\theta}(\omega)$ (Hinney et al. [2]). We also study the transmission of light that is detuned from the transition of the two-level emitter, a situation that we recently studied experimentally (Cordier et al. [3]). Our model predictions show how the collectively enhanced nonlinear response of weakly coupled emitters can be harnessed to generate non-classical states of light using ensembles ranging from a few to many emitters.
@@ I Introduction
Here, we treat the problem of many two-level emitters that are weakly coupled to a single propagating light mode. Our approach is based on the insight that photon-photon correlations in the light scattered by a single two-level emitter result from two-photon interference [29, 30, 31, 32]. By generalizing this concept to many two-level emitters, we introduce an intuitively accessible model, which allows us to calculate the quantum non-linear response of the ensemble of emitters and to derive analytical expressions for the second-order quantum coherence function and the squeezing spectrum of the output light. In order to check the validity of our model, we compare its predictions to those obtained in [33], where it has been shown theoretically that continuous wave input laser light can be transformed into perfectly anti-bunched or strongly bunched light depending on the number of coupled emitters, and find perfect agreement. Compared to previous theoretical descriptions, the simplicity of our approach results in a significant computational advantage and provides a much more intuitive picture. Furthermore, it can account for different probing configurations, e.g., through the waveguide or via external illumination, and different spatial arrangements of the emitters. Using our model predictions, we exemplify how to exploit the collective nonlinear response of an ensemble of emitters for, e.g., generating quantum light with tailored photon statistics or for realizing a narrowband source of squeezed light. Finally, while the transformation of laser light into antibunched light with a single quantum emitter is well-established [34] and was recently also shown to be possible with large ensembles of weakly coupled emitters [1], here, we anticipate that this is also possible with a small number (for example $N=3$) of suitably arranged emitters by illuminating them simultaneously form the outside and through the waveguide. This makes our model useful and attractive to a broad community working with different platforms, ranging from solid-state approaches, where the number of spectrally identical emitters is typically limited to a few [35], to quantum nanophotonics with trapped cold atoms, where the number of spectrally identical emitters can reach thousands.
@@
This article is organized as follows: In Sec. II, we introduce the fundamental quantity of this article, the two-photon wavefunction $\psi(\tau)$. We then decompose the two-photon wavefunction into its coherent and incoherent components and work out its relation to the second-order quantum correlation $g^{(2)}(\tau)$ function as well as to quadrature squeezing spectrum $S_{\theta}(\omega)$. In Sec. III we explain the significance of the coherent and incoherent component with the pedagogical example of resonance fluorescence from a single emitter where the resulting antibunched light can be understood as a destructive interference both components [29, 30, 31, 32]. Furthermore, we discuss the different illumination configuration for a single emitter (Sec. III.1 - Sec. III.3). In Sec. IV we introduce our model which extends the interference concept from a single emitter to an ensemble of emitters coupled to the same waveguide mode. We then apply our model to case studies under different excitation geometries. First, we recover with our model known results for emitters resonantly excited through the waveguide [33] (Sec. IV.1), and then extend it to new geometries such as illumination under the Bragg (Sec. IV.2) and anti-Bragg angle (Sec. IV.3), as well as a combined illumination from outside the waveguide and through the waveguide (Sec. IV.4).
@@ II Definitions
We seek to model the transformation of a weak classical continuous-wave monochromatic light field upon interaction with an ensemble of $N$ two-level emitters. We assume the latter to be weakly coupled to the same propagating optical mode, which is defined, e.g., by a single-mode optical waveguide  [^1] While, here, we limit the discussion to waveguide coupled emitters, the approach can be readily extended to emitters in free-space, such as in [65].. This situation can be experimentally realized with various emitters, such as laser cooled atoms, quantum dots, color centers, molecules in organic host material, or excitons, coupled to nanofibers [28, 37, 38], or to photonic waveguides [39, 40, 41, 42, 43, 35, 44, 45, 46, 47, 48, 49]. In the microwave regime, this can be also realized with microwave qubits coupled to a transmission line [50, 51]. While realizations with laser cooled atoms allow one to couple large numbers of weakly-coupled identical emitters [52, 18, 24, 25], solid state emitters allow up-to-now only the coupling of few identical emitters, with, however, much larger coupling strength [53, 54, 55, 56]. The model we present in this manuscript allows us to account for both, few and many weakly coupled identical emitters.
The aim of our model is to be able to describe the output quantum state of light in the waveguide when the ensemble of emitters is weakly driven with continuous-wave monochromatic laser light with frequency $\omega_{L}$, free-space wavelength $\lambda=2\pi c/\omega_{L}$, detuning $\Delta=\omega_{L}-\omega_{0}$ from the atomic transition frequency $\omega_{0}$, and the resonant Rabi frequency $\Omega$. The latter is complex, $\Omega=|\Omega|e^{i\phi}$, where the phase $\phi$ is defined by the phase of the driving field in absence of emitters. This general framework can describe different illumination geometries, through the waveguide, from the outside, or both, through the waveguide and from the outside, as illustrated in Fig. 1. Each of the $N$ emitters with population decay rate of the excited state $\Gamma$ is coupled to the forward-direction of a waveguide with a coupling efficiency $\beta=\Gamma_{\text{wg}}/\Gamma$ where $\Gamma_{\text{wg}}$ is the emission rate of the emitters into the forward-direction of the waveguide. In our model, we focus on the weak drive regime $\Omega\ll\Gamma$, where the driving field can be approximated by a truncated coherent state with at most two photons
@@
Using this truncation allows us to describe the output state after the interaction after the interaction with the ensemble of emitters as [33]:
@@
Figure 1: A single emitter coupled to a waveguide mode. The emitter can be driven externally ($\Omega^{\mathrm{ext}}$) and / or through the waveguide ($\Omega^{\mathrm{wg}}_{\mathrm{in}}$). Bottom: Possible one and two photons scattering processes.
The origin of the different components of the output state will become clearer when deriving them for distinct scattering processes. Each of the emitters can either coherently or incoherently scatter the incoming photons as illustrated in Fig. 1. As coherent scattering only involves a modification of the amplitude and phase, it is straightforward to account for. Calculating the impact of incoherent scattering on the output state turns out to be more involved. Incoherent scattering involves the scattering of two or more photons and, therefore, in our truncated Hilbert space, it only affects states with $X=2$. In the weak drive regime, where single-photon scattering dominates, the total output power is predominately governed by power of the coherent component of the output state, which scales linearly with the power of the driving field, $|\alpha_{\mathrm{out}}|^{2}\propto|\Omega|^{2}$. By contrast, the power of the incoherent component of the output state scales quadratically with the power of the driving field, $P_{\mathrm{incoh}}\propto|\Omega|^{4}$. However, while the incoherent component thus only contributes a negligible fraction to the total output power in the weak drive regime, it plays a crucial role in observables such as the normalized second-order correlation function, $g^{(2)}(\tau)$, or the squeezing spectrum, $S_{\theta}(\omega)$. These observables can be both inferred from the temporal two-photon wavefunction, $\psi(t_{1},t_{2})$, that is associated with |$\ket{\mathrm{out}}$. In the stationary regime considered here, it is given by:
@@
Here, the normalization $A$ is chosen such that the coincidence count rate is $G^{(2)}(\tau)=\langle\hat{a}^{\dagger}(0)\hat{a}^{\dagger}(\tau)\hat{a}(0)\hat{a}(\tau)\rangle=|\psi(\tau)|^{2}$ and $\hat{a}^{\dagger}(t)$ is the standard bosonic creation operator.
The two-photon wavefunction can be decomposed into the coherent and incoherent two-photon components $\psi(\tau)=\psi_{\mathrm{coh}}(\tau)+\psi_{\mathrm{incoh}}(\tau)$, cf. Fig. 1. In leading order of the driving field ($\Omega^{2}$), the normalized second-order correlation function is then given by [33]
$g^{(2)}(\tau)$ $=\frac{|\psi(\tau)|^{2}}{|\alpha_{\mathrm{out}}|^{4}}=\frac{|\psi_{\mathrm{coh}}(\tau)+\psi_{\mathrm{incoh}}(\tau)|^{2}}{|\alpha_{\mathrm{out}}|^{4}}.$ (4)
From Eq. (4), one can see that interference between the coherent and incoherent component of the two-photon wavefunction will impact the value of $g^{(2)}(\tau)$ [29, 58].
@@
Here, we introduced the generalized quadrature operator $\hat{X}_{\theta}(t)=\frac{1}{2}\left[\hat{a}_{t}e^{i\theta}+\hat{a}^{\dagger}_{t}e^{-i\theta}\right]$ with $\Delta\hat{X}_{\theta}(t)=\hat{X}_{\theta}(t)-\braket{\hat{X}_{\theta}(t)}$ where :::...:: … : denotes normal ordering. In the weak drive regime ($\Omega\ll\Gamma$), the squeezing spectrum is then approximately given by to [2, 60]:
@@
with $\varphi(\omega)=\text{Arg}\{\psi_{\mathrm{incoh}}(\omega)\}$. Equation (6) links the experimentally accessible squeezing spectrum $S_{\theta}(\omega)$ to the incoherent component of the two-photon wavefunction $\psi_{\mathrm{incoh}}(\omega)$, as demonstrated in [2].
@@ III Single emitter
Let us first consider the case of a single emitter coupled to a single mode waveguide as sketched in Fig. 1. The emitter is driven by a field that is either external or guided and has a resonant Rabi frequency $\Omega$ with $|\Omega|\ll\Gamma$ and a detuning $\Delta$. Correspondingly, in leading order the atom scatters photons coherently and incoherently with a total rate $R=|\Omega|^{2}/\Gamma(1+(2\Delta/\Gamma)^{2})$. The scattered photons are coupled into the forward direction of the guided mode with probability $\beta$, for which we assume the numerical value $\beta=0.01$ throughout the article, if not stated otherwise. The amplitude of coherently scattering a single photon into the forward direction of the guided mode is then
@@
The minus sign in Eq. (7) stems from the $\pi$ phase shift between the scattered light field and the driving field for on-resonant excitation  [^3] The $\pi$-phase shift on resonance is composed from two contribution: First, a $\pi/2$ from coupling into the waveguide (or the Gouy-phase in free-space) and, second, the dipole is oscillating with $\pi/2$ out-of-phase at resonance with respect to the excitation laser.. The amplitude of generating two coherently scattered photons in the waveguide is independent of their time difference $\tau$ and given by $\alpha_{\mathrm{sc}}^{2}$. The amplitude of incoherently scattering two-photons is given by [33, 62]:
@@
For $\tau=0$, the amplitude of this nonlinear scattering process is equal in magnitude to the coherent amplitude of scattering two photons into the waveguide $\psi_{\mathrm{coh}}$. For finite $\tau=t_{2}-t_{1}$, it exponentially decays with $|\tau|$. The photon-pairs created from an incoherent scattering are thus correlated in time with a typical separation $\tau_{\mathrm{exc}}=1/\Gamma$ – the lifetime of the excited state. This is in contrast with $\psi_{\mathrm{coh}}$ that describes two temporally uncorrelated photons.
In the frequency domain, the incoherent scattering can be understood as a degenerate four-wave mixing process where two laser photons are converted into a pair of red- and blue-detuned photons of frequency $\omega_{L}\pm\omega$ where $\omega$ is the detuning from the laser frequency $\omega_{L}$. The biphoton spectrum $\psi_{\mathrm{incoh}}(\omega)=\int_{-\infty}^{\infty}\psi_{\mathrm{incoh}}(\tau)e^{i\omega\tau}\mathrm{d}\tau$ is given by [62]:
@@ III.1 Single emitter, external illumination
In the case of external illumination of a single emitter with Rabi frequency $\Omega^{\mathrm{ext}}$, the coherent component of the two-photon wavefunction of the output guided field solely stems from the light scattered by the emitter, such that:
@@
Figure 3: Scattered two-photon wavefunctions from a single emitter at detuning $\Delta/\Gamma=1$: a) The coherent component $\psi_{\mathrm{coh}}$ is time-independent and has a constant phase. b) The incoherent component $\psi_{\mathrm{incoh}}(\tau)$ is a localized function, and the colors correspond to the phase difference to the coherent component $\Delta\varphi=\mathrm{Arg}\{\psi_{\mathrm{incoh}}(\tau)\}-\mathrm{Arg}\{\psi_{\mathrm{coh}}\}$. At $\tau=0$, the phase difference is $\pi$ such that in c) antibunching occurs. The second-order correlation function $g^{(2)}$ is deduced from a) & b) by taking into account the complex phase. The constructive/destructive nature of the interference is indicated by the background shading in b) & c), and its periodic change leads to Rabi-oscillation in $g^{(2)}(0)$.

### v2

# Present address: ]Istituto Nazionale di Ottica, Consiglio Nazionale delle Ricerche, 50019 Sesto Fiorentino, Italy thanks: These two authors contributed equally

@@ Abstract
We model and investigate the collective nonlinear optical response of an ensemble of two-level emitters that are weakly coupled to a single-mode waveguide. Our approach generalizes the insight that photon-photon correlations in the light scattered by a single two-level emitter result from two-photon interference to the case of many emitters. Using our model, we study different configurations for probing the nonlinear response of the ensemble, e.g., through the waveguide or via external illumination, and derive analytical expressions for the second-order quantum coherence function, $g^{(2)}(\tau)$, as well as for the squeezing spectrum of the output light in the waveguide, $S_{\theta}(\omega)$. For the transmission of resonant guided light, we recover the same predictions as previously made with more involved theoretical models when analyzing experimental results regarding $g^{(2)}(\tau)$ (Prasad et al. [1]) and $S_{\theta}(\omega)$ (Hinney et al. [2]). We also study the transmission of light that is detuned from the transition of the two-level emitter, a situation that we recently studied experimentally (Cordier et al. [3]). Our model predictions show how the collectively enhanced nonlinear response of weakly coupled emitters can be harnessed to generate non-classical states of light using ensembles ranging from a few to many emitters.
@@ I Introduction
Here, we treat the problem of many two-level emitters that are weakly coupled to a single propagating light mode. Our approach is based on the insight that photon-photon correlations in the light scattered by a single two-level emitter result from two-photon interference [29, 30, 31, 32, 33]. By generalizing this concept to many two-level emitters, we introduce an intuitively accessible model, which allows us to calculate the quantum non-linear response of the ensemble of emitters and to derive analytical expressions for the second-order quantum coherence function and the squeezing spectrum of the output light. In order to check the validity of our model, we compare its predictions to those obtained in [34], where it has been shown theoretically that continuous wave input laser light can be transformed into perfectly anti-bunched or strongly bunched light depending on the number of coupled emitters, and find perfect agreement. Compared to previous theoretical descriptions, the simplicity of our approach results in a significant computational advantage and provides a much more intuitive picture. Furthermore, it can account for different probing configurations, e.g., through the waveguide or via external illumination, and different spatial arrangements of the emitters. Using our model predictions, we exemplify how to exploit the collective nonlinear response of an ensemble of emitters for, e.g., generating quantum light with tailored photon statistics or for realizing a narrowband source of squeezed light. Finally, while the transformation of laser light into antibunched light with a single quantum emitter is well-established [35] and was recently also shown to be possible with large ensembles of weakly coupled emitters [1], here, we anticipate that this is also possible with a small number (for example $N=3$) of suitably arranged emitters by illuminating them simultaneously form the outside and through the waveguide. This makes our model useful and attractive to a broad community working with different platforms, ranging from solid-state approaches, where the number of spectrally identical emitters is typically limited to a few [36], to quantum nanophotonics with trapped cold atoms, where the number of spectrally identical emitters can reach thousands.
@@
This article is organized as follows: In Sec. II, we introduce the fundamental quantity underlying the physics of this article, the two-photon wavefunction $\psi(\tau)$. We then decompose the two-photon wavefunction into its coherent and incoherent components and work out its relation to the second-order quantum correlation $g^{(2)}(\tau)$ function as well as to the quadrature squeezing spectrum $S_{\theta}(\omega)$. In Sec. III we explain the significance of the coherent and incoherent component with the pedagogical example of resonance fluorescence from a single emitter where the resulting antibunched light can be understood as a destructive interference of both components [29, 30, 31, 37, 32, 33]. Furthermore, we discuss different illumination configurations for a single emitter coupled to a guided mode (Sec. III.1 - Sec. III.3). In Sec. IV, we introduce our model which extends the interference concept from a single emitter to an ensemble of emitters coupled to the same waveguide mode. We then apply our model to case studies of emitters coupled to a waveguide under different excitation geometries. First, we recover with our model known results for emitters resonantly excited through the waveguide [34] (Sec. IV.1), and then extend it to new geometries such as illumination under the Bragg (Sec. IV.2) and anti-Bragg angle (Sec. IV.3), as well as a combined illumination from outside the waveguide and through the waveguide (Sec. IV.4).
@@ II Definitions
We seek to model the transformation of a weak classical continuous-wave monochromatic light field upon interaction with an ensemble of $N$ two-level emitters. We assume the latter to be weakly coupled to the same propagating optical mode, which is defined, e.g., by a single-mode optical waveguide  [^1] While, here, we limit the discussion to waveguide coupled emitters, the approach can be readily extended to emitters in free-space, such as in [32].. This situation can be experimentally realized with various emitters, such as laser cooled atoms or molecules in organic host material coupled to nanofibers [28, 39, 40], or with quantum dots, color centers, or excitons coupled to photonic waveguides [41, 42, 43, 44, 45, 36, 46, 47, 48, 49, 50, 51]. In the microwave regime, this can be also realized with superconducting qubits coupled to a transmission line [52, 53]. While realizations with laser cooled atoms allow one to couple large numbers of weakly-coupled identical emitters [54, 18, 24, 25], solid state emitters allow up-to-now only the coupling of few identical emitters, with, however, much larger coupling strength [55, 56, 57, 58]. The model we present in this manuscript allows us to account for both, few and many weakly coupled identical emitters.
The aim of our model is to be able to describe the output quantum state of light in the waveguide mode when the ensemble of emitters is weakly driven with continuous-wave monochromatic laser light with frequency $\omega_{L}$, free-space wavelength $\lambda=2\pi c/\omega_{L}$, detuning $\Delta=\omega_{L}-\omega_{0}$ from the atomic transition frequency $\omega_{0}$, and the resonant Rabi frequency $\Omega$. The latter is complex, $\Omega=|\Omega|e^{i\phi}$, where the phase $\phi$ is defined by the phase of the driving field in the absence of emitters. In our model, we explore different illumination geometries for emitters coupled to a waveguide. The emitters can be illuminated either through the waveguide, externally, or simultaneously through both pathways, as illustrated in Fig. 1. Each of the $N$ emitters with the population decay rate of the excited state $\Gamma$ is coupled to the forward-direction of a waveguide with a coupling efficiency $\beta=\Gamma_{\text{wg}}/\Gamma$ where $\Gamma_{\text{wg}}$ is the emission rate of the emitters into the forward-direction of the waveguide. In our model, we focus on the weak drive regime $\Omega\ll\Gamma$, where the driving field can be approximated by a truncated coherent state with at most two photons
@@
Using this truncation allows us to describe the output state after the interaction with the ensemble of emitters as [34]:
@@
Figure 1: Top: A single emitter coupled to a waveguide mode. The emitter can be driven externally ($\Omega^{\mathrm{ext}}$) and / or through the waveguide ($\Omega^{\mathrm{wg}}_{\mathrm{in}}$). Bottom: Possible one and two photons scattering processes with the coherent single photon scattering amplitude $\alpha_{sc}$, the coherent two-photon scattering amplitude $\alpha_{sc}^{2}$ and the incoherent two-photon scattering amplitude $\psi_{\mathrm{incoh}}(\tau)$ which is given in Eq. (10).
The origin of the different components of the output state will become clearer when deriving them for distinct scattering processes. Each of the emitters can either coherently or incoherently scatter the incoming photons  [^3] Here, coherent and incoherent scatterign refers ti the first-order coherence properties of the two components, as commonly used in literature [89]. as illustrated in Fig. 1. As coherent scattering only involves a modification of the amplitude and phase, it is straightforward to account for. Calculating the impact of incoherent scattering on the output state turns out to be more involved. Incoherent scattering involves the scattering of two or more photons and, therefore, in our truncated Hilbert space, it only affects states with $X=2$. In the weak drive regime, where single-photon scattering dominates, the total output power is predominately governed by power of the coherent component of the output state, which scales linearly with the power of the driving field, $P_{\mathrm{coh}}\propto|\Omega|^{2}$. By contrast, the power of the incoherent component of the output state scales quadratically with the power of the driving field, $P_{\mathrm{incoh}}\propto|\Omega|^{4}$. However, while the incoherent component thus only contributes a negligible fraction to the total output power in the weak drive regime, it plays a crucial role in observables such as the normalized second-order correlation function, $g^{(2)}(\tau)$, or the squeezing spectrum, $S_{\theta}(\omega)$. These observables can be both inferred from the temporal two-photon wavefunction, $\psi(t_{1},t_{2})$, which in the stationary regime depends only on the relative time $\tau=t_{2}-t_{1}$, and is associated with |$\ket{\mathrm{out}}$ via the relation:
@@
Similarly, |$\ket{\mathrm{out}}$ can be represented in frequency space by $\psi(\omega)$ as
[…]
The two-photon wavefunction can be decomposed into the coherent and incoherent two-photon components $\psi(\tau)=\psi_{\mathrm{coh}}+\psi_{\mathrm{incoh}}(\tau)$, cf. Fig. 1, where the coherent component contains the coherently scattered component by the emitters, as well as - depending on the geometry - the residual of the driving field. For a continuous wave input the coherent component is time independent, while the incoherent two-photon component depends on the relative time $\tau$. In leading order of the driving field ($\Omega^{2}$), the normalized second-order correlation function is then given by [34]
$g^{(2)}(\tau)$ $=\frac{|\psi(\tau)|^{2}}{|\alpha_{\mathrm{out}}|^{4}}=\frac{|\psi_{\mathrm{coh}}+\psi_{\mathrm{incoh}}(\tau)|^{2}}{|\alpha_{\mathrm{out}}|^{4}}$ (5)
where $\alpha_{\mathrm{out}}$ is the coherent single photon component. From Eq. (5), one can see that interference between the coherent and incoherent component of the two-photon wavefunction will impact the value of $g^{(2)}(\tau)$ [29, 37].
@@
Here, we introduced the generalized quadrature operator $\hat{X}_{\theta}(t)=\frac{1}{2}\left[\hat{a}_{t}e^{i\theta}+\hat{a}^{\dagger}_{t}e^{-i\theta}\right]$ with $\Delta\hat{X}_{\theta}(t)=\hat{X}_{\theta}(t)-\braket{\hat{X}_{\theta}(t)}$ where :::...:: … : denotes normal ordering. In the weak drive regime ($\Omega\ll\Gamma$), the squeezing spectrum is then approximately given by [2, 62]:
@@
[…]
@@ III Single emitter
Let us first consider the case of a single emitter coupled to a single mode waveguide as sketched in Fig. 1. The emitter is driven by a field that is either external or guided and has a resonant Rabi frequency $\Omega$ with $|\Omega|\ll\Gamma$ and a detuning $\Delta$. Correspondingly, in leading order the atom scatters photons coherently and incoherently with a total rate $R=|\Omega|^{2}/\Gamma(1+(2\Delta/\Gamma)^{2})$. The scattered photons are coupled into the forward direction of the guided mode with probability $\beta$. The amplitude of coherently scattering a single photon into the forward direction of the guided mode is then
@@
The minus sign in Eq. (8) stems from the $\pi$ phase shift between the scattered light field and the driving field for on-resonant excitation  [^4] The $\pi$-phase shift on resonance is composed from two contribution: First, a $\pi/2$ from coupling into the waveguide (or the Gouy-phase in free-space) and, second, the dipole is oscillating with $\pi/2$ out-of-phase at resonance with respect to the excitation laser.. The amplitude of generating two coherently scattered photons in the waveguide is independent of their time difference $\tau$ and given by $\alpha_{\mathrm{sc}}^{2}$. The amplitude of the incoherently scattered two-photon component in forward direction is given by [34, 64]:
@@
For $\tau=0$, the amplitude of this nonlinear scattering process is equal in magnitude to the coherent amplitude of scattering two photons into the waveguide $\alpha_{\mathrm{sc}}^{2}$. For finite $\tau=t_{2}-t_{1}$, it exponentially decays with $|\tau|$. The photon-pairs created from an incoherent scattering are thus correlated in time with a typical separation $\tau_{\mathrm{exc}}=1/\Gamma$ – the lifetime of the excited state. This is in contrast with $\psi_{\mathrm{coh}}$ that describes two temporally uncorrelated photons.
In the frequency domain, the incoherent scattering can be understood as a degenerate four-wave mixing process where two laser photons are converted into a pair of red- and blue-detuned photons of frequency $\omega_{L}\pm\omega$ where $\omega$ is the detuning from the laser frequency $\omega_{L}$. The spectrum of the incoherent component is given by [64]:
@@ III.1 Single emitter, external illumination
In the case of external illumination of a single emitter with Rabi frequency $\Omega^{\mathrm{ext}}$, the coherent component of the two-photon wavefunction of the output guided field solely stems from the light scattered by the emitter, such that  [^5] Note that our definition follows [34, 62] which differs from [3, 32] by a factor $1/2$.:
@@
Figure 3: Scattered two-photon wavefunctions from a single emitter at detuning $\Delta/\Gamma=1$ undder external illumination: a) The coherent component $\psi_{\mathrm{coh}}$ is time-independent and has a constant phase. b) The incoherent component $\psi_{\mathrm{incoh}}(\tau)$ is a localized function, and the colors correspond to the phase difference to the coherent component $\Delta\varphi=\mathrm{Arg}\{\psi_{\mathrm{incoh}}(\tau)\}-\mathrm{Arg}\{\psi_{\mathrm{coh}}\}$. At $\tau=0$, the phase difference is $\pi$ such that in c) antibunching occurs. The second-order correlation function $g^{(2)}(\tau)$ is obtained from a) & b) by taking into account the complex phase. The constructive/destructive nature of the interference is indicated by the background shading in b) & c) such that at $\tau=0$ a fully destructive interference occurs. The interference pattern as function of $\tau$ then leads leads signatures of Rabi-oscillation in $g^{(2)}(\tau)$.